OpenStudy (anonymous):

What conditions need to be met for \[\lim _{x \rightarrow c} (\frac{g(x)}{f(x)})= \frac{\lim_{x \rightarrow c} g(x)}{\lim_{x \rightarrow c} f(x)}\] apart from the self-evident \[\lim_{x \rightarrow c} f(x) \ne 0\]?

OpenStudy (anonymous):

Are there any other restrictions?

OpenStudy (anonymous):

they both have to be continues at c

OpenStudy (anonymous):

example\[\lim_{x \rightarrow -1}\frac{\sqrt{x}}{x}=\frac{\sqrt{-1}}{-1}\] error

OpenStudy (anonymous):

oh wait not continuas but have a limit at -1

OpenStudy (anonymous):

or c

OpenStudy (anonymous):

Well, I was talking more generally about complex analysis, so that limit should be fine surely?

OpenStudy (anonymous):

Isn't that example just -i?

OpenStudy (anonymous):

OK, I'll take continuity, makes sense really. Thank you

OpenStudy (anonymous):

can we use the fact that the limit exist instead of continuity since a limit can exist even if the point is not continuous :shortly the \[\lim_{x \rightarrow c^+}f(x)=\lim_{x \rightarrow c^-}f(x)\]

OpenStudy (anonymous):

the lim has to exist for both functions